If f(x) =1/2x+1 and g(x)=2x-2, then f(g(-n))is
Answer by Mathtut(3670)

If f(x) =1/2x+1 and g(x)=2x-2, then f(g(-n))is.
:
g(-n)=2(-n)-2=-2n-2
:
f(g(-n))=(1/2)(-2n-2)+1=-n-1+1=-n